the length of a rectangle is 4 cm more than the width. the area is 96 cm^2. Find the dimensions of the rectangle?

Answer by checkley77(12844)

L=W+4
A=LW
96=(W+4)W
96=W^2+4W
W^2+4W-96=0
(W+12)(W-8)=0
W-8=0
W=8 ANS. FOR THE WIDTH.
L=8+4=12 ANS. FOR THE LENGTH.
PROOF:
96=12*8
96=96